# Robo-Instruct: Simulator-Augmented Instruction Alignment For Finetuning CodeLLMs

### v1

@@ 3 Analysis and Experiments
    Is Robo-Instruct effective at generating training data to fine-tune a small language model for generating domain-specific robot programs?
@@
We conduct our investigation by fine-tuning the Codellama-Python-7B model [30] on the synthetic dataset generated by Robo-Instruct and evaluate the fine-tuned model using RoboEval [10], a domain-specific code generation benchmark for service mobile robots. In the following subsections, we first provide a brief description of RoboEval. Then we present our experimental results addressing the two main research questions. Finally, we offer more analysis of RoboSim, InstAlign, and the synthetic dataset.
@@ 3.1 RoboEval: A Domain-Specific Robot Code Generation Benchmark
RoboEval is a domain-specific code generation benchmark, featuring a suite of 16 tasks designed to evaluate the ability of LLMs to understand custom APIs and generate programs for service robots. In this domain, a service robot can perceive objects, navigate to various locations, manipulate items, and communicate with humans. Furthermore, the robot should be capable of basic commonsense reasoning and executing complex tasks that involve conditional and repetitive actions. To facilitate these capabilities, RoboEval defines a set of 8 API functions in Python as skill primitives. Fig. 3 illustrates these function signatures and definitions, alongside an example task instruction and its canonical solution from the benchmark. In addition, unlike other popular code generation benchmark tasks [6, 2, 16, 19, 14, 9], the order of the robot’s actions is crucial for successfully completing the specified tasks. For instance, in the task "bring me a marker from the classroom that does not have a whiteboard," the robot must check each classroom until it finds one without a whiteboard, whereas simply bringing back a marker is insufficient. Hence, RoboEval evaluates the generated program by executing it in a simulator to capture the action traces, which are subsequently validated for sequence correctness using temporal logic.
@@
- GPT-3.5 - 67.5% 65.56% Proprietary
- Gemini-1.0-Pro - 60.00% 59.88% Proprietary
@@
Experiment Setup. We use the open-weight LLM, Llama3-8B-Inst, for Robo-Instruct. To generate a diverse dataset, we employ nucleus sampling for creating instruction-program pairs, setting the temperature $T=1$ and top $p=0.95$. The maximum resampling limit is capped at $3$ to accommodate instructions that initially produce invalid programs. For the LLM used in InstAlign, we empirically adjust the generation temperature to $T=0.3$ to optimize performance. Furthermore, we assess the edit similarity between token sequences of each instruction pair in the dataset [15], removing duplicates where the similarity score exceeds 0.6. We use the same setup to generate data via Self-Instruct. Instead of discarding invalid programs, Self-Instruct includes every generated instruction-program pair in the training dataset. Finally, we create two datasets with 5K instruction-program pairs each using Self-Instruct and Robo-Instruct respectively. These datasets are then used to fine-tune the Codellama-Python-7B model. The learning rate is set to be $-53e\text{-}$ 5 with a warmup ratio of $3\%$ and a constant lr scheduler. We employ the AdamW optimizer [20] with an effective batch size of 8, training each model for 5 epochs using a sequence length of 2048 tokens. We train all our models on a single H-100 GPU using unsloth [35].
@@
Baselines. We divide our baseline models into 2 categories: 1) proprietary LLMs, including GPT4 [28], GPT3.5-Turbo [27], Gemino-Pro [33], and 2) open-weight LLMs, including Codellama-Python-7B [30], Codellama-Python-34B, Starcoder2-33B [21], Deepseek-Coder-33B [8], and Llama3-8B-Inst [1]. All the results are evaluated using RoboEval and reported in Tab. 1.
@@
Tab. 1 presents the average pass@1 results for different LLMs on RoboEval, using two different temperature settings for generation: greedy decoding at a temperatures of $T=0$ and nucleus sampling at a temperature of $T=0.2$. The results show that Robo-Instruct-fine-tuned Codellama significantly improves upon the base Codellama-Python-7B and outperforms the Self-Instruct-fine-tuned variant. Notably, it surpasses all open-weight models, including larger ones like Codellama-Python-34B and Deepseek-Coder-33B. Additionally, although the training dataset was generated using Llama3-8B-Inst, which scores less than 50% pass@1 on RoboEval, our Robo-Instruct-fine-tuned model still achieves a significant improvement, scoring 68.75% under deterministic temperature settings for generation. Finally, compared to proprietary models, while our Robo-Instruct-fine-tuned model trails the more powerful GPT-4, it outperforms GPT-3.5-Turbo and Gemini-1.0-Pro in generating programs for service mobile robots. This result demonstrates the effectiveness of our approach in generating domain-specific robot program data for fine-tuning a small language model. It suggests that the fine-tuned model could potentially replace some proprietary models, providing a more cost-effective and private option for local deployment.
@@ 3.4 Qualitative analysis of the generated program errors
Figure 4: Self-Instruct-Generated Program Errors: Examples 1 to 4 illustrate errors specific to the Python language, and Examples 5 to 8 highlight errors rooted in domain-specific constraints. [^3] Programs have been adapted to succinctly demonstrate the types of errors and fit within the figure.
We analyze invalid programs identified by RoboSim, categorizing the errors into two types: language-native errors and domain-specific constraint violations. Fig. 3 displays eight examples of these programs, with Examples 1 to 4 illustrating errors specific to the Python language, and Examples 5 to 8 highlighting errors rooted in domain-specific constraints. Language-native errors are generally straightforward, such as syntax errors, the use of undefined variables or functions, or improper use of provided APIs.
In contrast, errors related to domain-specific constraints tend to be more complex to detect. For instance, Example 5 illustrates the program incorrectly trying to pick up a watering can (line 3) after establishing that it is not present at the location (line 2). Similarly, Example 6 demonstrates an error where the program inappropriately asks Jack (line 5) after confirming his absence from the room (line 3). Example 7 illustrates a scenario in which RoboSim updates the world state by labeling "item storage room" as a location after executing the go_to command (line 2). Subsequently, the robot attempts to pick up this location (line 3), resulting in an error. Example 9 is the most intricate scenario where the world state in the living room is updated to include a toy after the robot places it there (line 7). When the robot returns to the living room for the second time (line 5), it does not place down what it holds (line 7). Hence, in the third room the robot visits (line 3), when it attempts to pick up a toy again (line 4), an error occurs because the robot can only carry one item at a time.
@@ 4.1 LLMs for Robot Code Generation
LLMs have shown impressive capabilities in generating robot programs from natural language [17, 31, 11]. One popular approach uses LLMs to generate composable costmaps for robots to plan their motion on. In this approach, Voxposer [12] focuses on the tabletop manipulation setting and NavCon [3] focuses on creating composable maps for navigation. Using LLM to create reward functions is also promising. Eureka [23, 24] and Language to Rewards for Robotic Skill Synthesis [41] both show that LLM can generate good reward functions that allows robots to acquire complex skills. Finally, LLM can also be used to generate programs for high-level planning. LLM+p [18] outputs a robot plan in the form of the well-defined planning domain definition language (PDDL). Tidybot [39] uses an LLM to generate a rule that captures user preferences from examples and executes a program to sequentially complete the task in order. RoboEval [10] focuses on generating domain-specific programs for service mobile robots. It generates a program that allows the service robot to carry out long-horizon tasks and then validates the correctness of the program.
@@ 4.2 Generating Datasets For Fine-tuning LLMs
To enhance LLMs’ performance in code generation, numerous studies have explored the creation of specialized datasets [26, 13, 25]. Self-Instruct [36] is one popular method for generating synthetic datasets using an LLM. Following this methodology, Alpaca [32] generates 52K instruction-following demonstrations and subsequently fine-tunes the LLaMA 7B model [34] to create Alpaca 7B, which can behave qualitatively similarly to OpenAI’s text-davinci-003. Code Alpaca [5] extends this approach to generate code instructions using 21 seed tasks, while Gorilla-LM [29] adapts the method to focus on ML domain-specific APIs from Huggingface, TensorFlow Hub, and Torch Hub. To create more complex instructions, Evol-Instruct [40, 22] proposes iteratively updating instructions to become more complex through different prompting strategies. In addition to Evol-Instruct, OSS-Instruct [38] uses open-source code snippets to generate 75K high-quality instruction data and fine-tunes the Codelllama-Python-7B model to create Magicoder, which can match the performance of GPT-3.5-Turbo [27] on HumanEval [6]. While these works focus on creating seed instruction sets to generate synthetic data for effectively fine-tuning an LLM, our research investigates post-processing methods in addition to Self-Instruct. Specifically, we concentrate on generating domain-specific programs in robotics [10], where we can effectively leverage constraints to filter out erroneous programs.
@@ 5 Conclusion, Limitation and Future Works
In this work, we introduce Robo-Instruct, a novel framework to generate synthetic training data to fine-tune small language models for domain-specific robot programs. Robo-Instruct comprises two novel components: 1) RoboSim, an angelic-execution-based algorithm to effectively validate Self-Instruct-generated programs, and 2) InstAlign, an instruction alignment procedure to revise instructions to better align with the generated programs. The experimental results demonstrate that the Codellama-Python-7B model fine-tuned on the Robo-Instruct-generated dataset can significantly outperform many popular open-weight LLMs for generating domain-specific robot programs. It also outperforms two proprietary LLMs, GPT-3.5-Turbo and Gemino-1.0-Pro, as well as the Self-Instruct-fine-tuned variant. A limitation of this study is that Robo-Instruct relies on Self-Instruct to filter invalid programs, making the dataset quality dependent on Self-Instruct’s performance. This can introduce biases if Self-Instruct consistently fails in certain areas. Future work will explore integrating Robo-Instruct with advanced methods like Evol-Inst and OSS-Inst to enhance dataset quality for domain-specific robot programs.
@@ A.1 Overview
In this appendix, we first present ablation experiments to investigate the percentage of invalid programs generated by Self-Instruct and examine how the generation temperature in InstAlign affects final performance. Next, we analyze and compare the datasets generated by Robo-Instruct and Self-Instruct. Finally, we list the seed tasks used in RoboEvaland the CoT prompt.
A.2 Ablation Exmperiments
Figure 5: Ablation Experiments
A.2.1 effectivenss of the simulator
@@ A.2 Ablation Exmperiments
A.2.2 Instruction Alignment model temperature
We further investigate how varying LLM temperatures for generating $\mathcal{I}_{\text{revised}}$ in InstAlign impact the performance of the fine-tuned model. Fig. 5 shows the bar chart of the pass@1 score of the models fine-tuned over datasets generated using different LLM temperatures in InstAlign. The model performs the best when fine-tuned on the dataset generated using LLM temperature $T=0.3$. As the temperature increases, we observe a decrease in performance.
A.3 Analysis of the Generated Datasets
(a) Token Length Distribution for Self-Instruct vs.Robo-Instruct
(b) Cosine Similarity with RoboEvalfor Self-Instruct vs.Robo-Instruct
Figure 6: Dataset Analysis

### v2

@@ 3 Analysis and Experiments
    Is Robo-Instruct effective at generating training data to fine-tune a small language model to generate domain-specific programs for robots?
@@
We conduct our investigation by fine-tuning the CodeLlama-Python-7B model (Rozière et al., 2024) on the synthetic dataset generated by Robo-Instruct and evaluate the fine-tuned model using RoboEval (Hu et al., 2024), a domain-specific code generation benchmark for service mobile robots. In the following subsections, we first provide a brief description of RoboEval. Then we present our experimental results addressing the two main research questions. Finally, we offer more analysis of RoboSim, InstAlign, and the synthetic dataset.
@@ 3.1 RoboEval: A Domain-Specific Robot Code Generation Benchmark
RoboEval is a domain-specific code generation benchmark, featuring a suite of 16 tasks designed to evaluate the ability of LLMs to understand custom APIs and generate programs for service robots. In this domain, a service robot can perceive objects, navigate to various locations, manipulate items, and communicate with humans. Furthermore, the robot should be capable of basic commonsense reasoning and executing complex tasks that involve conditional and repetitive actions. To facilitate these capabilities, RoboEval defines a set of 8 API functions in Python as skill primitives. Fig. 5 illustrates these function signatures and definitions, alongside an example task instruction and its canonical solution from the benchmark. In addition, unlike other popular code generation benchmark tasks (Chen et al., 2021; Austin et al., 2021; Li et al., 2022; Liu et al., 2023b; Lai et al., 2022; Hendrycks et al., 2021), the order of the robot’s actions is crucial for successfully completing the specified tasks. For instance, in the task ”bring me a marker from the classroom that does not have a whiteboard,” the robot must check each classroom until it finds one without a whiteboard, whereas simply bringing back a marker is insufficient. Hence, RoboEval evaluates the generated program by executing it in a simulator to capture the action traces, which are subsequently validated for sequence correctness using temporal logic.
@@
- GPT-3.5-Turbo - 67.5% 65.56% Proprietary
- Gemini-Pro - 60.00% 59.88% Proprietary
@@
Experiment Setup. We use the open-weight LLM, Llama3-8B-Inst, for Robo-Instruct. To generate a diverse dataset, we employ nucleus sampling for creating instruction-program pairs, setting the temperature $T=1$ and top $p=0.95$. The maximum resampling limit is capped at $3$ to accommodate instructions that initially produce invalid programs. For the LLM used in InstAlign, we empirically adjust the generation temperature to $T=0.3$ to optimize performance. Furthermore, we assess the edit similarity between token sequences of each instruction pair in the dataset (Lee et al., 2022), removing duplicates where the similarity score exceeds 0.6. We use the same setup to generate data via Self-Instruct. Instead of discarding invalid programs, Self-Instruct includes every generated instruction-program pair in the training dataset. Finally, we create two datasets with 5K instruction-program pairs each using Self-Instruct and Robo-Instruct respectively. These datasets are then used to fine-tune the CodeLlama-Python-7B model. The learning rate is set to be $-53e\text{-}$ 5 with a warmup ratio of $3\%$ and a constant lr scheduler. We employ the AdamW optimizer (Loshchilov and Hutter, 2019) with an effective batch size of 8, training each model for 5 epochs using a sequence length of 2048 tokens. We train all our models on a single H-100 GPU using unsloth (Unslothai, 2024).
@@
Baselines. We divide our baseline models into 2 categories: 1) proprietary LLMs, including GPT4 (OpenAI et al., 2024), GPT3.5-Turbo (OpenAI, 2022), Gemino-Pro (Team et al., 2024), and 2) open-weight LLMs, including Codellama-Python-7B (Rozière et al., 2024), Codellama-Python-34B, Starcoder2-33B (Lozhkov et al., 2024), Deepseek-Coder-33B (Guo et al., 2024), and Llama3-8B-Inst (AI, 2024).
@@
Tab. 1 presents the average pass@1 results for different LLMs on RoboEval, using two settings: greedy decoding at temperature $T=0$ and nucleus sampling at temperature $T=0.2$. The results show that Robo-Instruct-fine-tuned CodeLlama significantly improves upon the base CodeLlama-Python-7B, and outperforms the Self-Instruct-fine-tuned variant (Appendix A.3.3 shows that improvements over Self-Instruct are not the result of distributional biases in the selection process). Notably, Robo-Instruct surpasses all open-weight models, including larger ones like CodeLlama-Python-34B and Deepseek-Coder-33B. Additionally, although the training dataset was generated using Llama3-8B-Inst, which scores less than 50% pass@1 on RoboEval from Tab. 1, our Robo-Instruct-fine-tuned model still achieves a significant improvement, scoring 68.75%. Finally, compared to proprietary models, while our Robo-Instruct-fine-tuned model trails the more powerful GPT-4, it outperforms GPT-3.5-Turbo and Gemini-Pro in generating programs for service mobile robots. This result demonstrates the effectiveness of our approach in generating domain-specific robot program data for fine-tuning a much smaller language model. It suggests that the fine-tuned model could potentially replace some proprietary models, providing a more cost-effective and private option for local deployment.
@@ 3.3 RQ2: How Do RoboSim and InstAlign Impact the Effectiveness of Robo-Instruct?
Figure 6: Self-Instruct-Generated Program Errors. Examples highlight errors that violate domain-specific constraints. [^4] Programs have been adapted to succinctly demonstrate the types of errors and fit within the figure.
@@ 3.4 Qualitative analysis of the generated program errors
We qualitatively analyze invalid programs identified by RoboSim, as shown in Fig. 4. The first three examples are easily recognizable to humans as flawed. However, the last example is more complex and involves an error when the robot can navigate to more than two rooms. After the robot places a toy in the living room, DynamicEval updates the environment to reflect that a toy is now in the room (line 8). However, when the robot returns to the living room later (line 6), it will not drop the item it’s holding (line 8). As a result, when the robot enters a third room (line 4) and tries to pick up another toy (line 5), an error will occur because the robot is only capable of carrying one item at a time. This example demonstrates that RoboSim can detect invalid programs beyond those easily identifiable through human inspection.
4 Real-World Deployment Results
We deployed the Robo-Instruct fine-tuned model to generate and execute mobile robot programs in the real world, as shown in Fig. 7. Compared to GPT-4 and GPT-3.5-turbo (Tab. 3), our model generates programs about 6x faster than GPT-4 and 3x faster than GPT-3.5-turbo, with similar output quality. In Appendix A.5 we showcase more results on long-horizon tasks beyond RoboEval.
Figure 7: Deployment of the Robo-Instruct fine-tuned model to generate programs based on user-provided instructions and execute them on the robot.
Models GPT-4 GPT-3.5-Turbo Robo-Instruct (Ours)
Inference Speed 19 tokens/s 42 tokens/s 114 tokens/s
Table 3: Inference speed of different models.
@@ 5.1 LLMs for Robot Code Generation
LLMs have shown impressive capabilities in generating robot programs from natural language (Liang et al., 2022; Singh et al., 2023; Huang et al., 2023a). One popular approach uses LLMs to generate composable costmaps for robots to plan their motion on. In this approach, Voxposer (Huang et al., 2023b) focuses on the tabletop manipulation setting and NavCon (Biggie et al., 2023) focuses on creating composable maps for navigation. Using LLM to create reward functions is also promising. Eureka (Ma et al., 2023, 2024) and Language to Rewards for Robotic Skill Synthesis (Yu et al., 2023) both show that LLM can generate good reward functions that allows robots to acquire complex skills. Finally, LLM can also be used to generate programs for high-level planning. LLM+p (Liu et al., 2023a) outputs a robot plan in the form of the well-defined planning domain definition language (PDDL). Tidybot (Wu et al., 2023) uses an LLM to generate a rule that captures user preferences from examples and executes a program to sequentially complete the task in order. RoboEval (Hu et al., 2024) focuses on generating domain-specific programs for service mobile robots. It generates a program that allows the service robot to carry out long-horizon tasks and then validates the correctness of the program.
@@ 5.2 Generating Datasets For Fine-tuning LLMs
To enhance LLMs’ performance in code generation, numerous studies have explored the creation of specialized datasets (Muennighoff et al., 2024; Köpf et al., 2023; Muennighoff et al., 2022). Self-Instruct (Wang et al., 2022) is one popular method for generating synthetic datasets using an LLM. Following this methodology, Alpaca (Taori et al., 2023) generates 52K instruction-following demonstrations and subsequently fine-tunes the LLaMA 7B model (Touvron et al., 2023) to create Alpaca 7B, which can behave qualitatively similarly to OpenAI’s text-davinci-003. Code Alpaca (Chaudhary, 2023) extends this approach to generate code instructions using 21 seed tasks, while Gorilla-LM (Patil et al., 2023) adapts the method to focus on ML domain-specific APIs from Huggingface, TensorFlow Hub, and Torch Hub. To create more complex instructions, Evol-Instruct (Xu et al., 2024; Luo et al., 2024) proposes iteratively updating instructions to become more complex through different prompting strategies. In addition to Evol-Instruct, OSS-Instruct (Wei et al., 2023) uses open-source code snippets to generate 75K high-quality instruction data and fine-tunes the CodeLlama-Python-7B model to create Magicoder, which can match the performance of GPT-3.5-Turbo (OpenAI, 2022) on HumanEval (Chen et al., 2021). While these works focus on creating seed instruction sets to generate synthetic data to effectively fine-tune an LLM, our research investigates post-processing methods in addition to Self-Instruct. Specifically, we concentrate on generating domain-specific programs in robotics (Hu et al., 2024), where we can effectively leverage constraints to filter out erroneous programs.
@@ 6 Conclusion, Limitation and Future Works
In this work, we introduce Robo-Instruct, a novel framework to generate synthetic training data to fine-tune small language models for domain-specific robot programs. Robo-Instruct comprises two novel components: 1) RoboSim, a method to synthesize consistent simulation environments dynamically for verifying the generated programs, and 2) InstAlign, an LLM-aided instruction alignment procedure to revise instructions to better align with the generated programs. The experimental results show that the 7B CodeLlama model fine-tuned on the Robo-Instruct dataset significantly outperforms larger open-weight LLMs and proprietary models like GPT-3.5-Turbo and Gemini-Pro in generating service robot programs. However, a key limitation is that Robo-Instruct relies on Self-Instruct to filter invalid programs, which may introduce biases and affect dataset quality. Another limitation is the use of a simple rejection-sampling method to handle invalid programs, which may not fully address the underlying issues. Future work will focus on improving dataset quality by integrating Robo-Instruct with advanced methods like Evol-Inst and OSS-Inst.
@@ A.1 Overview
In this appendix, we first outline the relationship between DynamicEval and the classic STRIPS planning formulation in subsection A.2, providing a new perspective on the proposed algorithm. In subsection A.3, we present additional ablation experiments to analyze the percentage of invalid programs generated by Self-Instruct and the effectiveness of the rejection-sampling strategy combined with RoboSim. We also explore how the generation temperature in InstAlign impacts final performance and compare the dataset diversity produced by Robo-Instruct and Self-Instruct. Subsection A.4 lists the seed tasks used in RoboEval and the CoT prompts. in subsection A.5, we report real-world experiments that empirically evaluate the performance of our fine-tuned model on two long-horizon tasks, which differ significantly from those in RoboEval, and assess the model’s latency in generating programs. Although this work focuses on service mobile robots, the proposed framework is adaptable to other domains. In subsection A.6, we offer toy examples showing how DynamicEval can be extended to verify programs by incorporating domain-specific constraints.
A.2 Relevance to STRIPS planning
The proposed DynamicEval shares significant similarities with the formulation of STRIPS planning. A STRIPS instance is typically represented as a tuple $\langle I,G,A,P\rangle$, where $I$ denotes the initial state of the simulation environment, $G$ represents the desired goal state that the robot aims to achieve, $A$ defines the set of actions available to transition between states, and $P$ is the set of preconditions that must be satisfied before performing actions. Thus, DynamicEval can be reformulated to align with the STRIPS formulation as shown in Alg. 3. Each API invocation corresponds to an action, and its precondition consists of a set of literals, representing specific combinations of entities, types, and states.
To address this, we extend the classic STRIPS formulation by incorporating dynamically discovered literals. Unlike the conventional STRIPS approach, where each literal is binary—True when defined and False when not—we introduce a third value, ”Undefined.” This means a literal must be explicitly defined as either True or False; otherwise, it remains in the Undefined state. When an action requires a literal that is undefined, a random value (True or False) is assigned to it, and the literal is added to the state of the simulation environment (line 7). Once the precondition is fully defined, the action is executed, and domain-specific constraints are checked for any violations (line 10). This extension enables DynamicEval to handle arbitrary programs effectively.
[…]
1:  Input: api_fn $\triangleright$ The API function name
2:  Input: api_inputs $\triangleright$ The input received by the API invocation
3:  Input: $\mathcal{W}$triangleright The current state of the simulation environment
[…]
5:  for $l\in p$ do $\triangleright$ Loop through every literal in the precondition
6:   if $\textsc{CheckDefined}(\mathcal{W},l)\textbf{ is }\text{Undefined}$ then
7:    $W\leftarrow\textsc{GrowWorld}(l,\mathcal{W})$triangleright Randonly instantiate the literal and grow $\mathcal{W}$ to include it
8:   end if
9:  end for
10:  $\text{retval},\mathcal{W}\leftarrow\textsc{ExecUpdate}(\text{api\_fn},\text{params},\mathcal{W})$triangleright Execute api_fn and update $\mathcal{W}$
11:  return  retval
A.3 Ablation Exmperiments
Figure 8: Ablation Experiment Results
A.3.1 the Effectiveness of the Rejection-Sampling Strategy
@@ A.3 Ablation Exmperiments
A.3.2 Instruction Alignment model temperature
We further investigate how varying LLM temperatures for generating the revised instruction in InstAlign impact the performance of the fine-tuned model. Fig. 8 shows the bar chart of the pass@1 score of the models fine-tuned over datasets generated using different LLM temperatures in InstAlign. The model performs the best when fine-tuned on the dataset generated using LLM temperature $T=0.3$. As the temperature increases, we observe a decrease in performance.
A.3.3 Analysis of Generated Dataset
Figure 9: Cosine similarities between RoboEval and generated data.
Figure 10: Token length distribution for each instruction-program pair.
Similar to Magicoder [Wei et al., 2023], we show the improvements from Robo-Instruct are not merely due to selection bias, i.e., including data more aligned with the distribution of RoboEval tasks than Self-Instruct. We pair each sample from the generated dataset with task instructions and their canonical solutions, then compute cosine similarity using TF-IDF embeddings [Sparck Jones, 1988]. Fig. 10 shows comparable token similarities between both methods. Additionally, Fig. 10 presents the token length distribution, which also appears similar for both.